Grade 11Mechanics

The magnitude of impulse developed by a mass of 0.2 kg which changes its velocity from 5î-3j+7k m/s to 2î+3j+k m/s

ashutosh verma
8 Years agoGrade 11

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear Ashutosh
 
As we know
 
Impulse = change in momentum
 = m * (V2 - V1)
 = 0.2 *(-3i + 6j -6k)
 For magnitude
 = 0.2 * sqrt(81) = 0.2 *9 
= 1.8 kg m/sec
 
Regards
Arun (askIITians forum expert)